A dc electric motor consumes 36000000 Joules when connected to a 0.25 kV supply for 60 minutes. Find the current taken from the supply?

Answer by choosing the closest answer of this question from the following choices.


a. 40 W

b. 30 A

c. 40 A

d. 10000 W


1
Expert's answer
2021-05-24T16:34:41-0400

Gives

Energy (E)=36000000J

Supply voltage (V)=0.25KV

time =60min =3600sec

We know that

"P=\frac{E}{t}\rightarrow(1)"

"P=IV\rightarrow(2)"

equation (1)and(2) equal

"\frac{E}{t}=IV"

"I=\frac{E}{tV}"

"I=\frac{36000000}{3600\times250}" =40A

Option (c) is correct option